OpenStudy (anonymous):

The curve y=ax^2+bx+c passes through points (1,8) (0,5) and (3,20) Find the values of a, b and c.

OpenStudy (anonymous):

start by plugging in (0,5) and see what you get

OpenStudy (anonymous):

5=c

OpenStudy (anonymous):

really good

OpenStudy (anonymous):

now plug in (1,8) using c=5 and see what you get

OpenStudy (anonymous):

8=a+b+5

OpenStudy (anonymous):

or 3=a+b now plug in (3,20) using c=5 and see what you get

OpenStudy (anonymous):

20=9a+3b+5

OpenStudy (anonymous):

or 15=9a+3b Now, you have two simultaneous linear eqs that you can solve by substitution or elimination to come up with a and b and you have then answered the question

OpenStudy (anonymous):

brilliant.

OpenStudy (anonymous):

how did you know how to solve this?

OpenStudy (anonymous):

I did not. I have been studying math for many years and always try to plug in a zero to see what happens. In this case it led me to find c and then since I had 2 other points, I started to see where it would lead.